OpenStudy (math2400):

can someone please help. A sucrose solution is prepared to a final concentration of 0.170M . Convert this value into terms of g/L, molality, and mass % (molecular weight, MWsucrose = 342.296g/mol ; density, ρsol′n = 1.02g/mL ; mass of water, mwat = 961.8g ). Note that the mass of solute is included in the density of the solution

OpenStudy (math2400):

first part is 54.76

OpenStudy (anonymous):

Can you describe the molality by definition?

OpenStudy (math2400):

like the actual definition? ...

OpenStudy (anonymous):

yes

OpenStudy (math2400):

one mol of solute/ kJ of solute ...

OpenStudy (anonymous):

great! can we now calculate the number of moles present in the sucrose?

OpenStudy (math2400):

umm do i have to use the molarity first..? idk cuz i don't have those value straight up..

OpenStudy (anonymous):

yes. we will have to figure out the number of moles present in the 0.170 M solution.

OpenStudy (anonymous):

we will have to calculate the number of moles of sucrose / the kilogram of the solvent.

OpenStudy (australopithecus):

The definition of molality is Mole solute/kg solvent it is used because volume is dependent on temperature therefore it is a concentration that is temperature independent. 1. pick an arbitrary volume of solution, I would go with 1L 2. convert mole to grams of sucrose, you can use this formula mol = grams/molecular mass 3. Now you have grams and volume of solution %(m/v) = grams of solute/Volume solution 4. You know how much water by mass in grams is in your solution convert it to kilograms, use moles of sucrose, and use the formula for molality I posted above these steps 5. you know grams of solute (sucrose) from solving %(m/v), you need to know total mass of solution for %(mass/mass), so you will need to use density of your solution. You also know you have 1L of solution (this is just arbitrary but you have to stick with it), convert 1L to mL then set up a ratio using the density to solve for the mass of the solution Knowing density of solution is: ρsol′n = 1.02g/mL 1.02g/1mL = x g/1 L %mass = mass of solute/mass of solution (both of the units of mass must cancel out, so mass of solute and mass of solution must both be in grams, or both in kg)

OpenStudy (australopithecus):

if you have any questions feel free to ask